Find the current in a long, straight wire needed to develop a magnetic field of    

       5x10-7 T at 6m from the wire.


1
Expert's answer
2022-01-25T07:11:12-0500

We shall find the current which is needed to develop the magnetic field of "5" X "10^{-7}T"


From the question, we have the parameters

"B=5" X "10^{-7}T"

length, "x=6m"

Current, "I=?"


Now,

Using the formula

"\displaystyle B=\frac{\mu_{o} I}{2\pi x}"


We shall make "I" the subject of the equation

We have


"\displaystyle I=\frac{2\pi x B}{\mu_{o}}"


where "\mu_{o}=4\pi \times 10^{-7}"


Substituting into the formula, we have

"\displaystyle I=\frac{2\pi \times 6\times 5\times10^{-7}}{4\pi \times 10^{-7}}"


"\displaystyle I=\frac{1 \times 30\times10^{-7}}{2 \times 10^{-7}}"


"\displaystyle I=\frac{30\times10^{-7}}{2 \times 10^{-7}}"


"\displaystyle I=\frac{30}{2}"


"I=15A"

which is the needed current to develop the magnetic field of "5\times10^{-7}T"